OpenStudy (anonymous):

parabola y=x^2+2x+L intersects line y=mx+n at points

OpenStudy (anonymous):

question continues here!! parabola y=x^2+2x+L intersects line y=mx+n at points \[\left( \alpha,\gamma1 \right) and \left( \beta,\gamma2 \right). if \alpha + \beta=3 and \alpha \beta=-4 \] then m+n is?

OpenStudy (vishweshshrimali5):

Okay intersection points are found out by solving the equations of both the curves. So, firs try to solve both the curves.

OpenStudy (vishweshshrimali5):

That is put y = mx +n in equation of parabola.

OpenStudy (anonymous):

i cant understand @vishweshshrimali5

OpenStudy (vishweshshrimali5):

Its okay

OpenStudy (vishweshshrimali5):

See: By solving the equations I mean, solve this: mx +n = x^2 + 2x + L

OpenStudy (vishweshshrimali5):

\[\large{x^2 + x(2-m) + (L-n) = 0}\]

OpenStudy (vishweshshrimali5):

Now since the solutions of x are \(\large{\alpha,\beta}\). Find out : (1) \(\large{\alpha+\beta}\) with the help of above equation. (2) \(\large{\alpha\beta}\) with the help of above equation.

OpenStudy (vishweshshrimali5):

Can you do this @dinisha ?

OpenStudy (anonymous):

is \[\alpha + \beta = 2-m\] and \[\alpha \beta = L-n\]

OpenStudy (vishweshshrimali5):

See: in \(\large{\alpha + \beta = \cfrac{-(2-m)}{1}}\) \(\large{\alpha\beta}\) is correct.

OpenStudy (vishweshshrimali5):

Remember that : sum of roots was -b/a .

OpenStudy (anonymous):

so -2 + m ?

OpenStudy (vishweshshrimali5):

Yes

OpenStudy (vishweshshrimali5):

But, you are already given \(\alpha+\beta,\alpha\beta\) in the question. So, with the help of those values find L and m.

OpenStudy (anonymous):

how??

OpenStudy (vishweshshrimali5):

See: in the question \(\large{\alpha+\beta = 3}\) But, we obtained that: \(\large{\alpha+\beta = -2 + m}\) So, -2+m = 3 => m = 5

OpenStudy (vishweshshrimali5):

Similarly, find out L

OpenStudy (anonymous):

here there are two unknows!! L-n=-4

OpenStudy (vishweshshrimali5):

Yeah sorry But, you now have one equation also.

OpenStudy (anonymous):

one equation??

OpenStudy (vishweshshrimali5):

I meant L-n = -4

OpenStudy (vishweshshrimali5):

Are you sure there was L in the original question ?

OpenStudy (vishweshshrimali5):

I think that's "1" not "L".

OpenStudy (vishweshshrimali5):

Now, the equation (L-n =-4) will become, 1- n = -4 => n = 5 So, m = 5 and n = 5. Now, can you find m+n ?

OpenStudy (vishweshshrimali5):

Did you get it @dinisha ?

OpenStudy (anonymous):

ohh yaa i got it thank you!!